# PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games

### v1

@@ 10 Do you have any personal opinions or theories about the case?
Detailed script statistics and evaluation question metrics are presented in TableA1.
B.2 Overall Performance Computing
@@ C.2 Comparison Models
Considering the strategic core of the game, a direct confession from the murderer would compromise its competitive nature. Hence, in all the methodologies we explore, we maintain that the murderer’s responses follow a predefined template, detailed in C.6.2. This method preserves the game’s integrity and level of challenge.
@@
    The SSA (Single Script Access)
    The SSA assesses agents’ performance when limited solely to their scripts, serving as a baseline for initial search efforts.
@@
    The FSA (Full Script Access)
    The FSA evaluates performance under conditions of unrestricted access to all scripts, representing an ideal search endpoint.
C.3 Performance Evaluation Results for English Dataset
@@
Script Evaluation #QA SSA FSA Agent’s Response After Playing the Game
@@
Table A4: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. SSA and FSA stand for Single Script Access and Full Script Access, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
@@ C.6.2 Gameplay Prompt
As outlined in sections 2.3, the Search by Questioning Prompt is developed based on data acquired from sensors. It includes variables like {victim}, {character}, {current_script}, {dialog_history}. Additionally, it integrates a summary, {summary}, synthesized from the sensor-collected data. The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the 3.3 chapter.
@@ C.6.3 Evaluation Prompt
Single-Choice Question Template for FSA Prompt
This template is specifically designed for Full Script Access (FSA), distinguishing it from conventional templates by granting access to the scripts of all players. Unlike standard procedures, it leverages the RAG technique to extract pertinent information from each player’s script for comprehensive analysis.
@@
Multiple-Choice Question Template for FSA Prompt
Similar to the single-choice version, this template is tailored for answering multiple-choice questions under FSA conditions.

### v2

@@ B.2 Evaluation questions
Detailed script statistics and evaluation question metrics are presented in Table1.
B.3 Overall Performance Computing
@@
B.4 Annotation Details
We recruited four human experts to read all the scripts and write the questions and answers. We conducted training sessions for them and organised game sessions to let them play and become familiar with the game flow. They were compensated at an hourly rate of \$31.92, with each narrative estimated to take about 5-20 hours to complete, depending on the complexity of the game.
B.5 Disclaimers of Risks
Consider that the detective narrative dataset may contain descriptions of violent events, actions, or characters, we warn annotators of such content before the annotation process:
Please be advised that the detective narrative dataset may contain descriptions of violent events, actions, or characters. This content is included solely for academic, research, and narrative analysis purposes. It is not meant to glorify or trivialise violence in any form. Annotators should be aware of potential exposure to violent content and engage with the dataset professionally and responsibly.
@@ C.2 Comparison Models
    The PP (Personal Perspective)
    The PP assesses agents’ performance when limited solely to their scripts, serving as a baseline for initial search efforts.
@@
    The OP (Omniscient Perspective)
    The OP evaluates performance under conditions of unrestricted access to all scripts, representing an ideal search endpoint.
C.3 Performance Evaluation Results
Script Evaluation #QA PP OP Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice PLAYER*
Death Wears White (9 players, 1 victim) Voting -- .$000_{\pm.000}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective $10$$233_{\pm.058}.$$467_{\pm.116}.$$333_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.058}.$$267_{\pm.116}.$
Ghost Revenge (7 players, 3 victims) Voting -- .$222_{\pm.416}.$$333_{\pm.471}.$$222_{\pm.416}.$$111_{\pm.314}.$$333_{\pm.471}.$$333_{\pm.471}.$
Objective $19$$333_{\pm.030}.$$334_{\pm.132}.$$368_{\pm.052}.$$316_{\pm.190}.$$404_{\pm.080}.$$509_{\pm.133}.$
Danshui Villa (7 players, 2 victims) Voting -- .$167_{\pm.373}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective $12$$389_{\pm.048}.$$222_{\pm.127}.$$167_{\pm.000}.$$305_{\pm.048}.$$306_{\pm.096}.$$361_{\pm.127}.$
Unfinished Love (7 players, 2 victims) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective $12$$167_{\pm.000}.$$167_{\pm.083}.$$167_{\pm.144}.$$278_{\pm.048}.$$111_{\pm.127}.$$222_{\pm.127}.$
Cruise Incident (5 players, 1 victim) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$333_{\pm.471}.$$333_{\pm.471}.$
Objective $4$$167_{\pm.289}.$$250_{\pm.250}.$$500_{\pm.250}.$$500_{\pm.000}.$$583_{\pm.382}.$$667_{\pm.144}.$
Sin (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$333_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$333_{\pm.471}.$
Objective $3$$000_{\pm.000}.$$333_{\pm.334}.$$111_{\pm.192}.$$111_{\pm.192}.$$111_{\pm.192}.$$333_{\pm.334}.$
Deadly Fountain (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$667_{\pm.471}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective $3$$000_{\pm.000}.$$556_{\pm.193}.$$000_{\pm.000}.$$111_{\pm.192}.$$000_{\pm.000}.$$111_{\pm.192}.$
Unbelievable Incident (5 players, 1 victim) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$$000_{\pm.000}.$
Objective $4$$083_{\pm.144}.$$083_{\pm.144}.$$333_{\pm.144}.$$167_{\pm.144}.$$083_{\pm.144}.$$250_{\pm.000}.$
Desperate Sunshine (4 players, 1 victim) Voting -- .$000_{\pm.000}.$$667_{\pm.471}.$$667_{\pm.471}.$$000_{\pm.000}.$$667_{\pm.471}.$ $1.000_{\pm.000}$
Objective $3$$333_{\pm.000}.$$556_{\pm.193}.$$556_{\pm.193}.$$333_{\pm.000}.$$667_{\pm.334}.$$778_{\pm.192}.$
Riverside Inn (4 players, 1 victim) Voting -- .$333_{\pm.471}.$ $1.000_{\pm.000}$$333_{\pm.471}.$$000_{\pm.000}.$ $1.000_{\pm.000}$$333_{\pm.471}.$
Objective $3$$444_{\pm.193}.$$889_{\pm.192}.$$444_{\pm.193}.$$000_{\pm.000}.$$667_{\pm.000}.$$556_{\pm.193}.$
Solitary Boat Firefly (6 players, 4 victims) Voting -- .$167_{\pm.373}.$$583_{\pm.493}.$$000_{\pm.000}.$$250_{\pm.433}.$$167_{\pm.373}.$$500_{\pm.500}.$
Objective $20$$267_{\pm.104}.$$500_{\pm.050}.$$183_{\pm.058}.$$350_{\pm.132}.$$217_{\pm.076}.$$517_{\pm.104}.$
Manna (6 players, 3 victims) Voting -- .$000_{\pm.000}.$$000_{\pm.000}.$$222_{\pm.416}.$$333_{\pm.471}.$$444_{\pm.497}.$$556_{\pm.497}.$
Objective $24$$250_{\pm.042}.$$389_{\pm.086}.$$403_{\pm.064}.$$444_{\pm.064}.$$430_{\pm.087}.$$569_{\pm.105}.$
Overall Voting -- .$074_{\pm.117}.$$326_{\pm.342}.$$176_{\pm.257}.$$058_{\pm.115}.$$245_{\pm.326}.$$282_{\pm.308}.$
Objective $117$$256_{\pm.121}.$$370_{\pm.184}.$$293_{\pm.157}.$$333_{\pm.140}.$$328_{\pm.187}.$$407_{\pm.227}.$
Table A3: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
For English Dataset
@@ C.3 Performance Evaluation Results
Script Evaluation #QA PP OP Agent’s Response After Playing the Game
@@
Table A4: Compare the performance of agents with other multi-agent algorithms designed for multiplayer deduction games. PP and OP stand for Personal Perspective and Omniscient Perspective, respectively, representing the performance of agents when they have access to either only their own script or the scripts of all agents, without interacting with other agents.
Each script’s complexity is indicated by the #$\#Tokens$ column, reflecting the narrative depth and the number of characters, suspects, and victims involved. This complexity directly influences the API usage cost, as more intricate scenarios require more processing power for information handling.
Script #Tokens Stage PP OP Agent’s Response After Playing the Game
Werewolf O-CoT ThinkThrice PLAYER*
Evaluation $0.349$ $1.433$ $0.815$ $0.807$ $0.790$ $0.799$
Evaluation $0.418$ $1.945$ $1.006$ $1.004$ $0.995$ $1.012$
Evaluation $0.351$ $1.415$ $0.834$ $0.816$ $0.833$ $0.825$
Evaluation $0.230$ $0.872$ $0.487$ $0.502$ $0.499$ $0.494$
Evaluation $0.064$ $0.327$ $0.162$ $0.165$ $0.164$ $0.163$
Sin (4 players, 1 victim) 2,121 Gameplay -- -- $0.680$ $0.833$ $0.730$ $0.603$
Evaluation $0.061$ $0.287$ $0.141$ $0.140$ $0.141$ $0.142$
Evaluation $0.045$ $0.207$ $0.107$ $0.111$ $0.111$ $0.109$
Evaluation $0.077$ $0.291$ $0.159$ $0.163$ $0.161$ $0.162$
Evaluation $0.104$ $0.383$ $0.222$ $0.221$ $0.224$ $0.220$
Evaluation $0.055$ $0.223$ $0.120$ $0.124$ $0.121$ $0.123$
Evaluation $0.380$ $1.571$ $0.816$ $0.811$ $0.814$ $0.823$
Evaluation $0.444$ $1.864$ $0.944$ $0.965$ $0.937$ $0.953$
Overall 47,906 Gameplay -- -- $38.329$ $46.162$ $40.464$ $33.702$
Evaluation $2.578$ $10.819$ $5.813$ $5.831$ $5.789$ $5.825$
Table A5: Compare the costs in US dollars(\$) of calling Openai API across multi-agent algorithms in MMGs setting, with Gameplay and Evaluation Stage. #$\#Tokens$ represent the average length of each character’s script. Costs are reported for one complete gameplay and one evaluation process for each script.
@@ C.6.2 Gameplay Prompt
As outlined in sections 2.3, the Search by Questioning Prompt is developed based on data acquired from sensors. It includes variables like {victim}, {character}, {current_script}, {dialog_history}. Additionally, it integrates a summary, {summary}, synthesized from the sensor-collected data. The element {question_number} denotes the total questions permitted, with a detailed discussion on the optimal number of questions presented in the 4.3 chapter.
@@ C.6.3 Evaluation Prompt
Single-Choice Question Template for OP Prompt
This template is specifically designed for Omniscient Perspective (OP), distinguishing it from conventional templates by granting access to the scripts of all players. Unlike standard procedures, it leverages the RAG technique to extract pertinent information from each player’s script for comprehensive analysis.
@@
Multiple-Choice Question Template for OP Prompt
Similar to the single-choice version, this template is tailored for answering multiple-choice questions under OP conditions.